### v2

# Pseudo-Bifurcations in Stochastic Non-Normal Systems

Abstract
We challenge the prevailing emphasis on bifurcations as the primary mechanism behind many abrupt surges and transient extremes in complex systems and propose an alternative more universally applicable explanation based on non-normal dynamics. We demonstrate that linear or approximately linear stochastic systems near a dynamical attractor exhibit transient repulsive dynamics – termed pseudo-bifurcations – when interacting components are sufficiently asymmetric and hierarchically organized, i.e., non-normal. These pseudo-bifurcations produce early-warning signals commonly linked to bifurcations, such as dimension reduction, critical slowing down, and increased variance. Furthermore, we show that, as actual bifurcations approach, non-normal transients also emerge, complicating their distinction and potentially creating a bias that suggests the system is much closer to a critical point than it actually is. We support our analytical derivations by empirically demonstrating that the brain exhibits clear signs of such non-normal transients during epileptic seizures. Many systems suspected of approaching critical bifurcation points should be reconsidered, as non-normal dynamics offer a more generic explanation for the observed phenomena across natural, physical, and social systems.
keywords:
non-normal system , transient dynamics , critical transitions , early-warning signals , tipping points
Many of the dynamics driving complex systems in nature and society are marked by abrupt changes that result in transitions to fundamentally different behaviors [1, 2, 3, 4]. In recent decades, the dominant framework for understanding these transitions has centered around the concept of bifurcations [5, 6, 4, 7]. A bifurcation is a point at which a small change in a system’s control parameter causes a qualitative shift in its equilibrium state, resulting in a new and distinct configuration of system behavior. There are numerous examples of systems that display dynamics resembling those of bifurcations. For instance in ecology, predator-prey dynamics can lead to the sudden extinctions of a species [8], small shifts in climate patterns may induce inter-glacial periods [9], vegetation losses above a critical threshold results in desertification [10], or nutrient overload can cause a lake to suddenly transition from a clear-water to a algae dominated state [11]. In social systems, notable examples are financial markets experiencing long periods of low volatility followed by abrupt crashes [12] and growing dissatisfaction with political conditions, which can trigger widespread unrest or revolution [13]. Health-related examples involve the sudden onset of conditions like microbiome dysregulation [14], transitions between different states representing distinct illnesses based on varying internal and external influences [15], depressive episodes [16], or epileptic seizures [17].
As a system approaches a critical bifurcation threshold, it is often marked by increased variance and autocorrelation in observable metrics. These statistical signatures reveal that the system’s recovery from perturbations is slowing as it approaches bifurcation, alongside an increasing sensitivity to fluctuations in the system’s state–two complementary indicators that signal the proximity to a critical transition. It has thus been suggested that these statistical signatures serve as important early-warning signals for the onset of a looming bifurcation point [5, 6, 4, 7].
Despite the insights offered by bifurcation theory, the framework itself encounters at least two major challenges. First, it relies on the existence of a well-defined control parameter and its closeness to a critical threshold. Second, it assumes the presence of a driving force that pushes the system toward or beyond this threshold. These assumptions are not only challenging to verify empirically but also conceptually restrictive and lack general applicability. It is, therefore, not surprising that the question of whether certain systems are undergoing critical bifurcations continues to be a subject of debate. For example, evidence is mixed regarding whether the onset of epilepsy is truly driven by critical transitions [18, 17]. Similarly, it remains uncertain whether components of the Earth’s system, such as the Amazon rainforest or the Atlantic Meridional Overturning Circulation (AMOC), are approaching bifurcation points [19].
In this work, we present an alternative, more general mechanism that relies solely on a stable equilibrium attractor with linear non-normal stochastic dynamics around it. We demonstrate that small perturbations around the equilibrium produce transient repulsions - we dub them pseudo bifurcations in this context - with similar statistical signatures to those observed near genuine bifurcations. Much like the effective dimension reduction seen during the onset of a bifurcation, we find that non-normal transients are confined to movements within a two-dimensional subspace. We further find that these transients are accompanied by statistical markers such as critical slowing down and increased autocorrelation. We empirically illustrate our findings by analyzing EEG signals that capture brain activity during epileptic seizures, revealing clear evidence of non-normal dynamics. We further demonstrate that the approach of actual bifurcations is accompanied by non-normal transients, making it difficult to distinguish between the two phenomena and potentially creating a bias that suggests the system is much closer to a critical point than it actually is. Finally, we show that the onset of non-normal transients can be anticipated by measuring the system’s alignment along its non-normal mode, and provide a numerical algorithm for stable identification of these non-normal modes. This opens a new path towards the anticipation and control of instabilities in complex systems.
The only requirement to observe such transients is that the system exhibits sufficient non-normality - meaning its components interact within asymmetric hierarchical structures, a feature common in many social and natural systems [20, 21]. This characteristic is also found in non-network systems, such as turbulent flows, where non-normal dynamics arise from the interplay of viscous and advective terms, which leads to asymmetric, non-orthogonal interactions among velocity perturbations [22]. Also, from a mathematical perspective, our findings are extremely generic. To observe pseudo-bifurcation, all one needs to assume is that the Jacobian of the linearized system, around its equilibrium point, forms a non-normal matrix. This is a highly general condition, as non-normal matrices form a measure-one set in the space of all matrices. Our findings thus suggest that many systems previously thought to be approaching criticality may need to be reconsidered, as they could instead be stable yet non-normal systems. More generally, our results offer a different perspective on how to interpret precursory signals in dynamical systems and highlight the role of non-normal dynamics in improving our understanding and management of a wide range of physical, natural, and social systems.
Reduced Non-Normal Dynamics
We consider a general $n$-dimensional system dynamics
$\dot{\mathbf{x}}=\mathbf{A}~{}\mathbf{x}+B~{}\boldsymbol{\eta}$ (1)
corresponding to the linearization of a general dynamics in the neighborhood of a stable attractor. The vector $\mathbf{x}$ quantifies the distance from the stable attractor. The matrix $\mathbf{A}$, which governs the system’s interactions, is an $n\times n$ matrix with eigenvalues $\lambda_{1},\ldots,\lambda_{n}$, all of which are assumed to have negative real parts, ensuring that the equilibrium point $\mathbf{x}=0$ remains stable. The $n$-dimensional noise term $\boldsymbol{\eta}$ is assumed to be normally distributed and $B$ quantifies the amplitude of the noise term. For simplicity, we take $B$ as scalar but reintroducing its tensorial nature does not change the main results presented below.
We further assume that the matrix $\mathbf{A}$ is non-normal, meaning $\mathbf{A}^{\dagger}\mathbf{A}\neq\mathbf{A}\mathbf{A}^{\dagger}$. Mathematically, this is equivalent to at least one pair of eigenvectors being non-orthogonal [22, 23]. Phenomenologically, this property implies that the dynamics in (1) can exhibit transient repulsive amplification, where the system may experience significant deviations from the stable equilibrium before eventually returning to it.
Qualitatively, the non-normality of $\mathbf{A}$ can be interpreted as representing interactions that are inherently directional and hierarchically structured [20, 21]. Many systems across the natural and social sciences are known to exhibit non-normality, examples including turbulence in fluid dynamics [22], perturbations in ecosystems [24, 25], non-Hermitian quantum mechanics [26], population dynamics [27], synchronization of optoelectronic oscillators [28], amplification of neural activities [29], chemical reactions [30], network synchronization [31, 32, 33], neuronal networks [34, 35] and financial markets [36, 37]. This should not be surprising since non-normal matrices - and by extension, non-normal dynamics - are the norm rather than the exception. In mathematical terms, non-normal matrices constitute a measure-one subset within the space of all matrices, making them overwhelmingly more prevalent. In contrast, normal matrices form a measure-zero subset, making them a special and less typical case. This prevalence implies that most real-world systems are inherently non-normal, with dynamics that often exhibit characteristic behaviors like transient growth and sensitivity to perturbations.
Figure 1: (top) Schematic representation of equation (1) in its reduced form (2). (bottom) Associated transient dynamics for both $||\mathbf{x}||$ as a function of time (main) and in the reduced $(n,r)$-plane (inset).
Denote by $\mathbf{P}$ the basis transformation matrix that diagonalizes $\mathbf{A}$, that is $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{-1}$ with $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$. Because $\mathbf{A}$ is non-normal, $\mathbf{P}$ is non-unitary, that is transforming $\mathbf{A}$ into diagonal form requires not just rotating but also scaling. Denote by $\mathbf{P}=\mathbf{U}\Sigma\mathbf{V}^{\dagger}$ the matrix $\mathbf{P}$’s singular value decomposition (SVD) and by $\mathbf{u}_{i}$ the $i$-th column vector of unitary matrix $\mathbf{U}$. Matrix $\mathbf{A}$ can be decomposed into pairwise projection operations $\mathbf{A}\sim\sum_{i,j=1}^{n}\frac{\sigma_{i}}{\sigma_{j}}\mathbf{u}_{i}\mathbf{u}_{j}^{\dagger}$ where $\sigma_{i}$ denotes the $i$-th singular value in $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{n})$. It is useful to consider the condition number $\kappa$ of matrix $\mathbf{A}$, defined as the ratio of the largest to the smallest singular values, $\kappa=\sigma_{1}/\sigma_{n}$, where we assume that the values are sorted without loss of generality. For a normal system, $\kappa=1$, which means that $\Sigma$ reduces to the identity matrix and the SVD of $\mathbf{P}$ reduces to $\mathbf{P}=\mathbf{U}\mathbf{V}^{\dagger}$, ensuring its unitarity and thus $\mathbf{A}^{\dagger}\mathbf{A}=\mathbf{A}\mathbf{A}^{\dagger}$. This suggests that the system’s degree of non-normality is well captured by the matrix $\mathbf{A}$’s condition number $\kappa$, as shown in the SM. Other measures of non-normality include the Kreiss constant or Henrici’s departure from normality [23].
In the remainder of this article, we assume, for simplicity, that the first $n-1$ eigenvectors are all (nearly) orthogonal to each other, with only the last one being non-orthogonal. With this assumption, $\sigma_{1}\approx\sigma_{2}\approx\ldots\approx\sigma_{n-1}\approx 1\gg\sigma_{n}=\kappa^{-1}$. This simplifies some of the subsequent analysis, but does not qualitatively change our results. As shown in the SM Appendix, so long as $\sigma_{n-1}\gg\sigma_{n}$ an assumption that is empirically supported across numerous sociotechnical and natural systems [38] - the solution of (1) exhibits non-normal transients. The SM Appendix further details that the dynamics of (1) can effectively be decomposed into two components: one along the non-normal mode $\hat{\mathbf{n}}\sim\mathbf{u}_{n}$ and one along its orthogonal reaction mode $\hat{\mathbf{r}}\sim\sum_{i=1}^{n-1}\mathbf{u}_{i}$, such that equation (1) is reduced to an effective two-dimensional system
$\dot{n}$ $=-\alpha_{n}n~{}+\kappa^{-1}r+B~{}\eta^{n}+\mathcal{O}\left(\kappa^{-1}\right)$ (2a)
$\dot{r}$ $=-\alpha_{r}r+~{}~{}~{}\kappa n~{}+B~{}\eta^{r}+\mathcal{O}\left(\kappa^{-1}\right)$ (2b)
where $n$ and $r$ denote the dynamics projected onto the non-normal and reaction mode $\hat{\mathbf{n}}$ and $\hat{\mathbf{r}}$, respectively. The coefficients $\alpha_{r,n}$ are constants of order one (Figure 1, top). The eigenvalues of system (2) are $\lambda_{\pm}=-\delta_{+}\pm\sqrt{1+\delta_{-}^{2}}$ with $\delta_{\pm}=(\alpha_{r}\pm\alpha_{n})/2$. The associated eigenvectors are given by $\mathbf{v}_{\pm}\sim(\mu_{\pm},\kappa)$ with $\mu_{\pm}\sim\mathcal{O}(1)$. The larger $\kappa$ becomes, the closer the eigenvectors are to being approximately collinear.
If the system’s degree of non-normality is sufficiently large, the dynamics of (1) become punctuated by transient repulsion. This can be understood from equations (2) by ignoring the stochastic term and higher-order terms. Starting at $(n,r)=(1,0)$, the variable $n$, and thus $||\mathbf{x}||_{2}$, initially decreases as the non-normal mode relaxes to zero at a rate $\alpha_{n}$ via (2a). Simultaneously, the reaction component $r$ grows rapidly at a rate $\kappa$ via (2b). When $\kappa$ is sufficiently large, to be specified below, it dominates the relaxation dynamics, leading to the growth of $||\mathbf{x}||_{2}$. This growth saturates at its peak when $\dot{r}=0$, after which the relaxation dynamics takes over, driving the system back to its stable equilibrium at $||\mathbf{x}||_{2}=0$ at rate $\lambda_{+}$ (Figure 1, bottom). We stress that this two-dimensional decomposition applies to systems of arbitrary dimension $n$.
Characterization and Prediction of Non-Normal Transients
Figure 2: (left) Behavior of the Ornstein-Uhlenbeck parameter $\theta_{t}$ from (3) for different levels of non-normality $\kappa$. For $\kappa>\kappa_{c}$, the typically positive, mean-reversion coefficient $\theta_{t}$ transiently becomes negative, thereby inducing repulsive growth. (middle) Regime diagram as a function of the leading eigenvalue $\lambda_{+}$ of reduced system (2) and condition number $\kappa$. As the bifurcation at $\lambda_{+}=0$ is approached at fixed $\kappa$, the system first traverses the pseudo-critical regime as $\kappa_{c}$ becomes smaller than $\kappa$. (right) Schematic representation of the two regimes that dictate non-normal dynamics. Small values of $||\mathbf{x}||$ are driven by the stochastic component, but once the system is significantly aligned along the non-normal component $n$, quasi-deterministic, stochastic loops are induced.
To characterize the parameter regime in which the non-normal transients dominate the dynamics, it is helpful to rewrite equation (2b) as an effective Ornstein-Uhlenbeck process (or AR(1) process in discrete time)
$\dot{r}=-\theta_{t}~{}r+B~{}\eta^{r}$ (3)
where $\theta_{t}$ is a time-dependent mean-reversion rate with memory. The exact expression of $\theta_{t}$ is given in the SM and its qualitative behavior is effectively illustrated in Figure 2 (left), capturing the key properties derived in the appendix. For $\kappa>\kappa_{c}$, where
$\kappa_{c}\approx\sqrt{1+4\lambda_{+}\lambda_{-}}~{},$ (4)
$\theta_{t}$ becomes negative over a finite time interval, corresponding to a repulsion of $r$ from its stable equilibrium at $r=0$. For systems with $\kappa>\kappa_{c}$ the stochastic dynamics of $||\mathbf{x}||_{2}$ is characterized by intermittent shifts between two regimes. For relatively small values of $||\mathbf{x}||$, the noise is dominating and the behavior appears primarily stochastic around the equilibrium point. If, by chance, the system’s trajectory is increasingly parallel to the non-normal mode, the transient amplification takes hold and the system is dominated by the approximately deterministic transient repulsion that gives rise to stochastic loops (Figure 2, right).
We verify these results by simulating an $n=10$ dimensional system with $B=1$. The matrix $\mathbf{A}$ has negative eigenvalues $\lambda_{1}=-10,\lambda_{2}=-9,\ldots,\lambda_{10}=-1$, which determines $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{10})$. We further assume that the first $n-1$ singular values $\sigma_{1},\ldots,\sigma_{n-1}$ of transformation matrix $\mathbf{P}$ are equal to $1$. This allows us to tune the system’s degree of non-normality by setting the last singular value $\sigma_{10}$ equal to $\kappa^{-1}$. The matrix $\mathbf{A}$ is then generated via $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{-1}$ where $\mathbf{P}=\mathbf{U}\Sigma\mathbf{V}^{\dagger}$ with $\mathbf{U}$ and $\mathbf{V}^{\dagger}$ two randomly generated orthonormal matrices and $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{10})$. Figure 3 shows two dynamics, one where $\mathbf{A}$ is normal $(\kappa=1)$ and one where it is non-normal $(\kappa=30)$, with otherwise equal conditions. The system’s mean-field state value, $\langle\mathbf{x}\rangle\equiv\frac{1}{n}\sum_{i=1}^{n}x_{i}$, is punctuated by large repulsive transients compared to its normal counterpart. The quasi-deterministic long-memory transient behavior along the reaction mode $r$ is reflected in the complementary cumulative distribution of $||\mathbf{x}||$ that exhibits a mixture of two well separated contributions, one for normal fluctuations and one associated with large stochastic loops.
Figure 3: (a-c) Mean-field, effective eigenvalue and variance measured via (5) from a simulated system (1) with $n=10$ dimensions: normal case (grey line, $\kappa=1$) and non-normal case (black line, $\kappa=30$) with otherwise equal conditions. (e-g) Same as (a-c) but only for the non-normal system for a specific excerpt for better resolution. The vertical dashed line highlights a pseudo-bifurcation at which the effective eigenvalue crosses zero. (d) Reduced dynamics (2) in the $(n,r)$-plane. While the non-normal system (yellow to red) displays distinct transient repulsion and contraction along the reaction mode $r$, the normal mode (green to blue) shows no such trends (as also highlighted in the re-scaled inset figure). (h) Complementary cumulative distribution (CCD) of projection along the reaction mode $r$ from subplot (d). The CCD of the non-normal system (dark line) is a mixture of two distributions: one dominated by noise, and the other dominated by quasi-deterministic stochastic loops. The second distribution is absent in the normal case (grey line). The probability distribution function (PDF) of the mean-field (inset) shows no such distinction. (i) Autocorrelation of the mean-field dynamics. (j) Lead-lag correlation between the system’s alignment along the non-normal mode and the mean-field dynamics, with a clear indication that the former leads the later when the system is non-normal.
Assuming full knowledge of $\mathbf{A}$ an assumption that is relaxed below - further allows us to anticipate the onset of a non-normal transient. A predictor is the projection $n$ of the system onto the non-normal mode $\hat{\mathbf{n}}$, which dictates the transient growth along $\hat{\mathbf{r}}$ via (2b). The lead-lag correlation between $n$ and $\langle\mathbf{x}\rangle$ as a function of the lag $\tau$ shows that $n$ significantly leads the onset of large transients (Figure 3 (j)) in non-normal systems, with no such predictability in the normal case. This suggests that, by monitoring a system’s alignment along the non-normal mode $n$, one can anticipate the onset of a transient perturbation.
Pseudo-Bifurcations
A bifurcation in a continuous dynamical system occurs when a small change in a parameter causes a qualitative change in the stability of an equilibrium or fixed point. Mathematically, this corresponds to the real part of an eigenvalue of the equilibrium passing through zero. Formally, for a system $\dot{\mathbf{x}}=f(\mathbf{x};\mu)$, a bifurcation occurs at $(\mathbf{x}_{0},\mu_{0})$ if the Jacobian $Df_{\mathbf{x}_{0},\mu_{0}}$ has eigenvalues with zero real part. If the eigenvalue has an imaginary component, the bifurcation is classified as a Hopf bifurcation, leading to periodic limit cycles. Otherwise, it is referred to as a steady-state bifurcation, such as a saddle-node bifurcation (where two equilibria collide) or a pitchfork bifurcation (where one fixed point branches into three). These bifurcations are not just mathematical constructs but have significant real-world implications, leading to qualitatively altered system behavior. For instance, desertification in semi-arid regions, such as the Sahel, occurs when vegetation loss passes a critical threshold, leading to a rapid and often irreversible shift to desert landscapes [10]. Similarly, lake eutrophication, such as in Lake Erie, involves a bifurcation where nutrient overload causes a sudden transition from a clear-water state to a turbid, algae-dominated system [11].
Assume that we do not have perfect observation of $\mathbf{x}$, nor a clear understanding of the underlying dynamics, but that we observe the system’s mean-field state value $\langle\mathbf{x}\rangle$ (or similarly, its norm $||\mathbf{x}||$). We then model the system’s mean-field dynamics as an effective Ornstein-Uhlenbeck process
$\dot{\langle\mathbf{x}\rangle}=\phi~{}\langle\mathbf{x}\rangle+B_{\text{eff}}~{}\eta$ (5)
where $\phi\equiv\exp(\lambda_{\text{eff}})-1<0$ captures the system’s mean-reversal component with effective negative eigenvalue $\lambda_{\text{eff}}$. The system in equation (5) exhibits fluctuations with a variance equal to $B^{2}/(2|\phi|)$, and the correlation between states separated by a time interval $\Delta t$ is given by $\exp(-|\phi|\Delta t)$. Therefore, near a critical point, autocorrelation decays slowly with a characteristic time scale diverging as $\sim|\phi|^{-1}$, causing the system to relax slowly to equilibrium, a phenomenon known as critical slowing down. At the same time, the variance tends to infinity. These characteristic statistical signatures are collectively known as early-warning signals for bifurcation transitions [5, 6, 4, 7].
However, while bifurcations can generate such patterns, similar dynamics can also emerge when a non-normal system experiences transient repulsion, even in the absence of an actual bifurcation. Figure 3 shows the temporal evolution of $\phi$ and $\lambda_{\text{eff}}$ obtained by fitting an Ornstein-Uhlenbeck process to the mean-field dynamics (1). During large transients, these coefficients approach zero, while the autocorrelation and variance increase in a way similar to the onset of a critical bifurcation. These observations can be explained by the fact that the transient dynamics temporarily induce a regime of repulsive, quasi-deterministic growth via (2b). At a qualitative level, the stochastic loops (Figure 2, right and Figure 3, (d)) could easily be mistaken for Hopf limit cycles. This underscores that non-normality, much like bifurcations, can generate the same early-warning signals, complicating the interpretation of these indicators. Given the similarity between these statistical patterns and those observed near critical transitions, even when the underlying non-normal system is far from or entirely devoid of true critical behavior, we refer to these transient repulsive dynamics as pseudo-bifurcations. This terminology is further justified by the fact that the natural tool to analyse the transient dynamics of non-normal operators is the pseudospectrum [23], which extends the concept of eigenvalues by replacing the condition of a vanishing characteristic determinant with the requirement that it be small, thereby capturing transient amplification effects that standard spectral analysis misses.
Distinguishing true bifurcations (if any) from pseudo-bifurcations becomes even more challenging as a system nears an actual critical bifurcation point, because, as we now show, the likelihood of observing non-normal transients increases simultaneously.
Equation (4) shows that, as the system’s effective eigenvalue $\lambda_{\text{eff}}$, and thus $\lambda_{+}\lambda_{-}$, approaches zero, the critical threshold $\kappa_{c}$ at some point crosses the system’s non-normality level $\kappa$, and hence repulsive transients emerge stochastically. In other words, before entering a critical regime, the system first passes through a pseudo-critical regime (Figure 2, middle). Early warning signals that are usually used to gauge a system’s distance to criticality indicate wrongly that the system is much closer to a critical point than it actually is.
We emphasize that pseudo-bifurcations are observed at a fixed level of non-normality, $\kappa>\kappa_{c}$, within a stationary system (1) fluctuating around an equilibrium point. This provides a highly general first-order approximation of equilibrium dynamics without requiring the system’s control parameter to move toward a critical point. Given the prevalence of interdependence structures $\mathbf{A}$ known to exhibit non-normality [21], we propose that pseudo-bifurcations are likely far more common than true bifurcations in empirical systems.
In addition, complex systems are frequently governed by mechanisms such as proportional growth that drive $\mathbf{A}$ towards states of increased hierarchical organizations and hence non-normality [39, 21, 37]. The SM derives that the variance of system (5) scales like $B_{\text{eff}}\approx B\kappa^{2}/\left(\lambda_{+}\lambda_{-}(\lambda_{+}+\lambda_{-})\right)$. Similarly, the auto-correlation at lag $\Delta t$ is given by $\exp(-\Gamma(\tau)~{}\Delta t)$ with $\Gamma(\tau)\approx 1/B_{\text{eff}}$. As the system’s non-normality, and thus $\kappa$ increases, so does the system’s effective variance and auto-correlation time. This suggests an increased likelihood of false-positive markers of bifurcation in systems that dynamically self-organize into hierarchical structures.
Non-Normal Transients during Epileptic Seizure
Figure 4: (Top Left) Mean-field state of the 23 EEG channels before, during (red), and after an epileptic seizure. Inset plots show a high-resolution version of a single second before (blue) and during (red) the seizure. (Bottom Left) Condition number $\kappa$ of the EEG connectome estimated over a rolling window of 40 seconds (light grey line). An additional moving average over a time scale of 10 seconds is applied, giving the smoother black line. (Top Right) Stochastic loops, similar to Figure 3(d), along the non-normal and reaction components during an excerpt of non-seizure (left/blue) and seizure (right/red). (Bottom Right) Lead-lag correlation between the system’s exposure to the non-normal component and the mean-field state during a non-seizure (blue) and a seizure (red) period. The purple vertical line indicates when the correlation reaches its maximum. During the seizure, excitation along the non-normal component leads the mean-field excursions by about 0.1 seconds.
Brain connectomes - comprehensive maps of structural and functional neural connections - are essential tools for studying the brain’s complex dynamics. In the context of epilepsy, brain connectomes have been used to explore the neural underpinnings of seizures, which involve sudden, excessive synchronization of neural activity [40]. It has been hypothesized that seizures may be linked to the brain’s connectome operating near a critical state, a threshold where the brain’s dynamics are poised between order and chaos [41]. However, whether or not the brain operates at criticality during epileptic seizures remains debated [18, 17]. A mean-field model of the excitation and inhibition dynamics has been proposed to explain these phenomena [42]. This model demonstrates how non-normal transients, driven by an interplay between excitation and inhibition, can lead to large bursts of neural activity, reminiscent of seizure dynamics. Building on these observations, and because it has been shown that the connectome is strongly non-normal [21], we hypothesize that a non-normal transient burst in the brain’s dynamics may be the underlying source of the apparent increase in excitability during seizures, shifting the system into a more excitable, uncontrolled regime and revealing the inherent non-normal characteristics of the connectome.
To test this claim, we utilize an EEG recording of a patient during an epileptic seizure recorded at the Children’s Hospital Boston [43]. The measurement includes data across 23 EEG channels sampled at 256 Hz, before, during and after a seizure event that has been classified as lasting for 40 seconds (see top-left in Figure 4 for the mean-field dynamics). We first estimate the matrix $\mathbf{A}$ of (1) as $\hat{\mathbf{A}}$ obtained by fitting a VAR(1) process to the data over a rolling window of 40 seconds. In principle, we could now proceed by first determining the transition matrix $\mathbf{P}$ that diagonalizes $\hat{\mathbf{A}}$, and subsequently calculating the SVD $\mathbf{P}=\mathbf{U}\Sigma\mathbf{V}^{\dagger}$ to identify the non-normal mode $\hat{\mathbf{n}}\sim\mathbf{u}_{n}$ and reaction mode $\hat{\mathbf{r}}\sim\sum_{i=1}^{n-1}\mathbf{u}_{i}$. However, as we detail in the SM, this approach is highly sensitive to noise, especially when $\mathbf{A}$ is non-normal. Therefore, we propose an alternative approach that consists in estimating the non-normal and reaction component directly from $\hat{\mathbf{A}}$ as
$\hat{\mathbf{r}},\hat{\mathbf{n}}=\operatorname*{arg\,max}_{\mathbf{v},\mathbf{w}\in\mathbb{S}^{n-1}}\langle\mathbf{v},\hat{\mathbf{A}}\mathbf{w}\rangle~{}~{}\text{s.t.}~{}~{}\langle\mathbf{v},\mathbf{w}\rangle=0.$ (6)
where $\mathbb{S}^{n-1}$ is the unit sphere in $n$ dimensions. $\langle\mathbf{v},\hat{\mathbf{A}}\mathbf{w}\rangle$ represents the inner product (or dot product) between the vector $\mathbf{v}$ and the result of applying the matrix $\hat{\mathbf{A}}$ to the vector $\mathbf{w}$. It thus measures the degree of alignment or projection of the transformed vector $\hat{\mathbf{A}}\mathbf{w}$ onto $\mathbf{v}$. Equation (6) directly captures the intuition behind the reduced equation (2) whereby the non-normal component maximally projects onto the reaction component. From (2), we further see that $\kappa$ can be estimated directly from $\hat{\mathbf{r}}$ and $\hat{\mathbf{n}}$ via $\kappa=\sqrt{\left|\langle\hat{\mathbf{r}},\hat{\mathbf{A}}\hat{\mathbf{n}}\rangle\right/\left|\langle\hat{\mathbf{n}},\hat{\mathbf{A}}\hat{\mathbf{r}}\rangle\right|}$ thus allowing us to completely bypass the need for estimating the singular values which are numerically unstable. In the SM, we test our empirical procedure on synthetic data, finding that our procedure faithfully recovers the non-normal and reaction mode, and estimates $\kappa$ more precisely than when calculating it directly via singular value decomposition.
The lower-left plot of Figure 4 reveals a marked increase in the connectome’s non-normality that occurs precisely at the onset of the epileptic seizure, persisting well beyond its clinical definition. This quantitatively demonstrates that the dynamics associated with the seizure extend far beyond the observable clinical symptoms. This is further supported by the prolonged high amplitude and variance in EEG recordings, which remain elevated long after the seizure has clinically ended.
In line with the observation that $\mathbf{A}$ becomes highly non-normal after the onset of the epileptic seizure, we observe in the top-right plots of Figure 4 that there are stochastic loops during, but not before the seizure.
By measuring the lead-lag correlation between the system’s projection along $\hat{\mathbf{e}}_{n}$ and the mean-field state $\langle\mathbf{x}\rangle$, we observe that the former leads the later by about $0.1$ seconds (Figure 4, bottom-right). This suggests that epileptic seizures may not only be driven by non-normal dynamics but also that the onset of transient episodes could be anticipated by measuring the system’s exposure to non-normal components. This insight offers a promising pathway for enhancing the early detection and potential control of seizure events, while also paving the way for managing self-induced instabilities in complex systems more broadly [44].
Reexamining Criticality
Stanislaw Ulam famously quipped that “using a term like nonlinear science is like referring to the bulk of zoology as the study of non-elephant animals,” humorously highlighting that linear systems are rare in nature, while nonlinear dynamics dominate real-world phenomena. This remark underscores how the emphasis on linearity has often overshadowed the vastly more prevalent and complex behaviors of nonlinear systems. We propose that a similar oversight has occurred with respect to non-normal dynamics. In the realm of complex systems, most interaction matrices are inherently non-normal, meaning that transient dynamics driven by non-normality are not just common but should be anticipated as central features. In fact, the situation for non-normal systems is even more universal, as these transient behaviors arise even within linear non-normal systems. Just as linear systems have historically received disproportionate attention at the expense of nonlinear ones, bifurcations have similarly overshadowed the far more general and widespread influence of non-normal transients. We emphasize again that the dynamics of an arbitrarily large, complex system around an equilibrium point is well approximated, to first order, by (1), and that the Jacobian matrix $\mathbf{A}$ being non-normal is both mathematically and phenomenologically the norm, not the exception. Ironically, non-normal matrices are much more “normal” (i.e. common) than normal matrices.
Bifurcation theory has been instrumental in understanding critical transitions, offering profound insights into how systems behave as they approach or cross critical thresholds. However, it addresses only one specific scenario. In contrast, non-normal transients can occur under much broader conditions, even when a system is far from a bifurcation point. For instance, for (5) to exhibit critical behavior, the real part of the system’s largest eigenvalue has to be extremely close to 00, whereas non-normal transients can be observed for values well below this threshold. This suggests that many abrupt surges and transient extremes presumed to be critical might can, in reality, be better explained by non-normal transients rather than proximity to bifurcation points. Consequently, we believe such cases merit re-examination to uncover the true underlying dynamics. For instance, recent research has demonstrated that financial bubbles can be interpreted as sub-critical transient excursions driven by hierarchically organized influence networks [36] and that these networks self-organize towards non-normal configurations [37].
Naturally, we must apply our framework with caution. It is well-established that many systems experience true bifurcations, where crossing a critical threshold leads to a fundamental change in behavior. Such transitions result in a qualitative shift in dynamics, which is both observable and empirically validated. Beyond the cases of desertification [10] and lake eutrophication [11] previously mentioned, coral reefs present a notable example: under stressors like pollution, overfishing, and climate change, reefs can cross a threshold, transitioning from coral-dominated to algae-dominated states. This shift is often irreversible, as the algae-dominated state hinders coral regrowth through altered ecological dynamics [45]. Similarly, certain fisheries, such as the Newfoundland cod fishery, have experienced abrupt collapses due to overfishing; once population levels dropped below a critical threshold, recovery became impossible, disrupting both ecological and economic systems [46].
However, in many cases, the assumption that a system is nearing a critical point remains largely hypothetical, with empirical evidence still inconclusive. For instance, there is ongoing debate over whether epileptic seizures constitute a critical transition in brain dynamics. While some studies suggest that the brain operates near a critical point, others find the evidence inconclusive [18, 17]. Our findings align with the latter perspective, reinforcing the view that seizures do not necessarily correspond to critical transitions.
Similarly, the Atlantic Meridional Overturning Circulation (AMOC) is suspected of approaching a bifurcation point that could lead to abrupt climate shifts, yet definitive observations remain lacking [19]. It is also possible that warmer interglacial periods observed in Earth’s climate history do not result from crossing a critical bifurcation point but instead reflect non-normal transients within an otherwise glacial climate regime. In these examples, the system’s proximity to a critical point is speculative, and a re-evaluation through the lens of non-normal dynamics could offer new insights.
Conclusions
By focusing on non-normal dynamics, we have proposed an alternative explanation for early-warning signals traditionally attributed to bifurcations in complex systems. We demonstrated that systems with a high degree of non-normality can exhibit transient repulsive dynamics – called pseudo-bifurcations – which give rise to the same statistical signatures as true bifurcations, such as critical slowing down and increased variance. Importantly, these patterns do not necessarily indicate the proximity to a critical bifurcation point. Our empirical analysis of EEG recordings during epileptic seizures further supports this hypothesis, showing that the brain’s non-normal dynamics are prominently dominating epileptic seizures. These findings suggest that many systems previously believed to be approaching criticality may, in fact, be at long-term equilibrium but exhibit strong non-normal transients. This calls for a reexamination of systems previously surmised to operate near criticality and offers a new perspective on the anticipation and control of instabilities in dynamical complex systems.
Acknowledgements
This work is partially supported by the National Natural Science Foundation of China (Grant No. T2350710802 and No. 72350410487), Shenzhen Science and Technology Innovation Commission Project (Grant No. GJHZ202107-05141805017 and No. K23405006), and the Center for Computational Science and Engineering at Southern University of Science and Technology.